OpenStudy (anonymous):

dy/dx=(y-x)(4x+y)/x(5x-y)

OpenStudy (amistre64):

let y = vx y' = v + v'x

OpenStudy (anonymous):

Help me to solve this

OpenStudy (amistre64):

i just dd ...

OpenStudy (amistre64):

wherever you see a y, replace it by vx wherever you see a y', replace it by v+v'x it then becomes seperable

OpenStudy (amistre64):

or so ive been lead to assume :)

OpenStudy (anonymous):

then i should used dy=vdx+xdv?

OpenStudy (anonymous):

What would be the equation looked like? I cannot really understand how. So sorry

OpenStudy (amistre64):

dy/dx=(y-x)(4x+y)/x(5x-y) \[y'=\frac{(y-x)(4x+y)}{x(5x-y)}\] \[y=vx,~y'=v+v'x\] \[v+v'x=\frac{(vx-x)(4x+vx)}{x(5x-vx)}\] \[v+v'x=\frac{xx(v-1)(4+v)}{xx(5-v)}\] \[\frac{dv}{dx}x=\frac{(v-1)(4+v)}{(5-v)}-v\] can you see how its going?

OpenStudy (anonymous):

why didn't you used the dy=vdx+xdv?

OpenStudy (anonymous):

hey? still there?

OpenStudy (anonymous):

??

OpenStudy (amistre64):

i have my own notation i perfer

OpenStudy (amistre64):

\[y=vx\] \[\frac{d}{dx}y=\frac{d}{dx}vx\] \[\frac{dy}{dx}=\frac{dv}{dx}x+v\frac{dx}{dx}\] since dx/dx = 1 \[\frac{dy}{dx}=\frac{dv}{dx}x+v\] \[dy=x~{dv}+v~dx\] is fine

OpenStudy (amistre64):

its just easier to write it in ' notation is all

OpenStudy (amistre64):

\[v'x=\frac{(v-1)(4+v)}{(5-v)}-v\] \[v'x=\frac{(v-1)(4+v)-v(5-v)}{(5-v)}\] \[\int\frac{(5-v)}{(v-1)(4+v)-v(5-v)}~dv=\int\frac 1xdx\] pretty sure they constructed it so that left side simplifies

OpenStudy (amistre64):

\[\frac12\int\frac{5-v}{(v-2)(v+1)}dv=ln(x)+C\] looks like a simple decomp

OpenStudy (anonymous):

i just copy your answers but i need to prefer with my notes. i need to study this, our exam is on thurs. thank you! differential equation and mechanics is not easy for me

OpenStudy (anonymous):

if you prefer in my notes. does it look like this vdx+xdv/dx=(vx-x)(4x+vx)/x(5x-vx) ?

OpenStudy (amistre64):

thats doable yes

OpenStudy (anonymous):

then hows next?

OpenStudy (amistre64):

factor your xs ... you want to algebra this into 2 seperate function: f(v) and g(x)

OpenStudy (anonymous):

can ou show me how?

OpenStudy (amistre64):

i already did ....

OpenStudy (amistre64):

if you are working on diffy qs, you should already know some basic algebra to manipulate stuff around ... i got it too:\[\int\frac{(5-v)}{(v-1)(4+v)-v(5-v)}~dv=\int\frac 1xdx\] which can further be worked (assuming i didnt make an error) to:\[\frac12\int\frac{5-v}{(v-2)(v+1)}dv=\int \frac1x dx\]

OpenStudy (anonymous):

used my preference? can you?

OpenStudy (amistre64):

you can use your preferences ... either way, it will algebra into that to work on

OpenStudy (anonymous):

Thank you. Goodnight maybe i should get slomme sleep. till next time =))

OpenStudy (anonymous):

Take care

OpenStudy (amistre64):

sweet dreams :)